OpenStudy (anonymous):

square root of x-3=5

Nnesha (nnesha):

 i think question is

OpenStudy (anonymous):

\[\sqrt{x-3} = 5\] \[(\sqrt {x-3})^2 = 5^2\]

Nnesha (nnesha):

@jdog92644 ?????

OpenStudy (anonymous):

Wait. Is it \[\sqrt{x - 3} = 5\] or \[\sqrt{x} - 3 = 5\]

OpenStudy (anonymous):

\[x-3 = 25\] \[x = 25 +3\] x = 28